OpenStudy (anonymous):

Anybody help me prove this identity step-by-step, reason-by-reason please? Will give medal and fan. Prove the identity by simplifying both sides of the equation. (sin^2 x) * (csc^2 x) = sin^2 x + cos^2 x

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

You could have just not answered then? lol

OpenStudy (anonymous):

I will give you medal if you give me one?

OpenStudy (anonymous):

No thanks lol I seriously need help with this problem, I'm not tryina play games right now.

OpenStudy (anonymous):

I just wanted a medal but I will give you one because I am nice bye bye

OpenStudy (anonymous):

Alright, thanks.

hero (hero):

Actually, start with the LHS: \((\sin^2x)(\csc^2x)\) Convert \(\csc^2x\) to \(\dfrac{1}{\sin^2x}\): \((\sin^2x)\left(\dfrac{1}{\sin^2x}\right)\)

hero (hero):

Notice that \(1 = \sin^2x + \cos^2x\): \(\dfrac{\sin^2x}{\sin^2x}(\sin^2x + \cos^2x)\) And \(\dfrac{\sin^2x}{\sin^2x}\) cancels leaving just: \(\sin^2x + \cos^2x\)